Observer: E. Sani, M. Fumana
Telescope Operator: S. Anderson, D. Gonzales-Huerta
Support Astronomer: D. Thompson
Instrument: LBC+LUCI

February 14-15

Cirrus above the orizon (SW) at beginning of the night, with low wind and humidity. The initial seeing is 1.1 on the GCS.
During the first 4hrs both humidity and temperature go up and down, with some difficulties of the primary mirrors to thermalize.
Very thin cirrus passing.
During the night the average seeing is highly variable. See the attached plot as an example.
Good sky flats for the LBc Bin are taken during night 20130217.

Project_RUN Objects Filter ExpTime Notes
Standard SA98 B-Bessel, R-Bessel, U-Bessel, V-Bessel   airmass 1.46
Saracco DDT Linx B-Bessel, R-Bessel 2h the seeing is out of the given specs
Saracco DDT Linx U-Bessel, V-Bessel 2h 20% of the images are expected with seeing close to the specs
Standard SA101 B-Bessel, R-Bessel, U-Bessel, V-Bessel, U-special, r-Sloan   airmass 1.2
Grazian 30 GOODS-N U-special, r-Sloan 2h30m some frames are out of specs for the seeing
Standard SA104 U-special, r-Sloan   airmass 1.5

February 15-16

At the beginning of night operations there are cirrus scattered over the South sky, with wind gusts at 12 m/s and low humidity.
At the end of the twilight there remain cirrus above the SW orizon, but the rest of the sky is almost clear. The wind is increasing with gusts at 20 m/s.
After 3h from the beginning of the observation the sky is clear.
During the night the average seeing experience huge variations, ranging between 0.8 and 2.
With LUCI we have defocusing during the MOS acquisition. Its origin is currently unknown, but, considering that the effect is small and appears only in the upper part of the mask combined with the bad seeing, it should not affect seriously the scientific frames.
Good sky flats for the LBC Bin are taken during night 20130217.

Project_RUN Objects Filter ExpTime Notes
Standard SA98 B-Bessel, R-Bessel, U-Bessel, V-Bessel   airmass 1.4
Saracco DDT Linx B-Bessel, R-Bessel 1h  
Saracco DDT Linx U-Bessel, V-Bessel 1h  
Standard SA101 B-Bessel, R-Bessel, U-Bessel, V-Bessel    
Standard Hip 33836 HKspec,200_H+K    
Zamorani_11 zCgr32 HKspec,200_H+K 3h 30m 50% of the frames are within the seeing specs
Standard SA101 U-special, r-Sloan   airmass 1.2
Grazian 30 GOODS-N U-special, r-Sloan 1h in both channels sky and seeing conditions are within the specs

February 16-17

The sky is CLEAR at the beginning of the night, with very low humidity and low wind.
At the beginning of scientific integration the seeing on the DIMM is around 1.0.
The sky remains clear all night long.
During the night the seeing is around 1.0 or better.
LBC Bin sky flats taken during this night are good.

Project_RUN Objects Filter ExpTime Notes
Nascimbeni_DDT GJ3470b U-special, FN971N20 3h35m COMPLETE
Standard SA101 B-Bessel, R-Bessel, U-Bessel, V-Bessel   airmas 1.2
Saracco_DDT Lynx U-Bessel, V-Bessel 1h  
Standard SA104 U-special, r-Sloan    
Grazian_30A GOODS-N U-special, r-Sloan 2h sky conditions and seeing within the specs

February 17-18

Thin cirrus above the South orizion at the beginning of the night, with low humidity and quite strong wing gusts ~15m/s.
Due to the high wind gusts at 4.00UT we have to close the dome.
The sky is clear when we open again the dome.
During the execution of the Grazian program we had an trackers error from the red side. The TO fixed it and the SA re-initialized the software.

Project_RUN Objects Filter ExpTime Notes
Standard SA101 B-Bessel, R-Bessel, U-Bessel, V-Bessel   airmas 1.2
Saracco_DDT Lynx B-Bessel, R-Bessel 1h COMPLETE
Standard SA104 U-special, r-Sloan   airmass 1.6
Grazian_30A GOODS-N U-special, r-Sloan 2h sky conditions and seeing within the specs
Standard SA101 U-special, r-Sloan   airmass 1.6
Standard SA109 U-special, r-Sloan   airmass 1.4

February 18-19

During the late evening the sky is covered by very thick clouds. We execute LUCI calibrations
Problem with the red side door of the dome. Mike tries to open it manually from level 10
It is not possible to fix the problem, we observe with the blue-side alone

Project_RUN Objects Filter ExpTime Notes
Antoniucci_DDT HH34 FeII ~24min Interrupted: target close to wind dir
Zamorani_A1 Hip33836 HKspec    
Zamorani_A1 zCgr32 HKspec 2.5h Interrupted: guide star lost
LaFranca_A9 Hip061493 200HK/zJspec    
LaFranca_A9 NGC4338 200HK/zJspec 45min COMPLETE
Grazian_30A GOODS-N U-special 1h12min Only blue channel
Standard SA109 U-special    

February 19-20

Dome closed at the begining of the night because of strong wind and dewpoint limit.
At 02:00UT it is snowing.
We have problems to insert the calibration unit in LUCI. We call the SA: a power lost during the day caused the lack of energy to the calibration unit. The SA fix it.
Error setting the MOS instrument. We call Dave. A slit mask got stuck, and the SA fix the problem.
The Dome remains closed all night long.

Project_RUN Objects Filter ExpTime Notes

February 20-21

We stay close due to the snow.

Project_RUN Objects Filter ExpTime Notes

-- EleanoraSani - 20 Feb 2013
Topic revision: r2 - 28 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback